OpenStudy (anonymous):

Solve equation. Check for extraneous solution. 2|7-7x|=2x+4

OpenStudy (anonymous):

i know that 2|7-7x|=2x+4 and 2|-7+7x|=2x+4 but then what do i do

OpenStudy (anonymous):

First step should always be to isolate the absolute value, not to negate it. So: 2|7-7x|=2x+4 |7-7x|=x+2 THEN you split the left side into the negative and positive situations: -(7-7x)=x+2 AND (7-7x)=x+2 Then solve and check for extraneous solutions.

OpenStudy (anonymous):

how do you know if it is an extraneous solution or not

OpenStudy (anonymous):

You solve for x and then plug it into the original equation to see if it's true or not.

OpenStudy (anonymous):

your wrong the multiple choice question was a. x=3/2 or x=5/6 b. x=3/2 or x=5/8 c.x=5/8 d.x=3/2 the answers i got were x=2/3 or x=8/5

OpenStudy (anonymous):

someone else have better help??

OpenStudy (primeralph):

Yeah @Ell . Show your skills. @karli111 She's good at this stuff.

OpenStudy (primeralph):

@Ell We discussed similar problems earlier and recently.

OpenStudy (primeralph):

Well, I just wanted to have you both learn in tandem.

OpenStudy (primeralph):

@karli111 Here goes:

OpenStudy (primeralph):

Nobody's a pro by default. Takes practice courage. Forget about it. Good luck.

OpenStudy (primeralph):

@karli111 Still there?

OpenStudy (anonymous):

yeah why

OpenStudy (primeralph):

Okay, I'll explain it. Pay attention:

OpenStudy (primeralph):